# Higher-Order Staircase Codes

### v1

@@ IV-A Code Structure
Consider starting with a generalized staircase code of sidelength $S^{\prime}$ and memory parameter $M^{\prime}$. We further have our permutations of $[S^{\prime}]\times[S^{\prime}]$ given by $\pi^{\prime}_{k}$ for $k\in[M^{\prime}+1]$ and the corresponding matrix-permuting functions given by $\Pi^{\prime}_{k}$ for $k\in[M^{\prime}+1]$. Finally, we have a ruler $d_{0}<d_{1}<\dots<d_{M^{\prime}}$ and a component code $\mathcal{C}$ of length $(M^{\prime}+1)S^{\prime}$ and dimension $(M^{\prime}+1)S^{\prime}-r$. However, we have the following twist: We define our constraint that the rows of
@@
belong to $\mathcal{C}$ as holding for all integers $n\in L\mathbb{Z}$ where $L$ is a new fixed parameter. In other words, we restrict the constraint on the rows of (14) to integers $n$ that are multiples of $L$.
Revisiting the Tanner graph perspective, the constraint nodes become $\mathcal{L}=L\mathbb{Z}\times[S]$ and from (5) we see that a variable node $(n,i,j)\in\mathcal{P}$ has the constraint $(n^{\prime},i^{\prime})\in\mathcal{L}$ as a neighbour if and only if $n^{\prime}=n+d_{k}$ for some $k\in[M^{\prime}+1]$ where $n^{\prime}\in L\mathbb{Z}$. Therefore, we get a variable node degree of
@@
which depends on the block index $n$ and the ruler. We seek to recover our original variable degree of $M+1$ for all variable nodes independently of $n$. This motivates the following definition:
@@
An $L$-uniform ruler of order $M^{\prime}+1$ is a set of integers $d_{0}<d_{1}<\dots<d_{M^{\prime}}$ for which
|$\absolutevalue{\{d_{k}\mid k\in[M^{\prime}+1],d_{k}\equiv\ell\pmod{*}{L}\}}=\frac{M^{\prime}+1}{L}$
for each $\ell\in[L]$. Necessarily, $M^{\prime}+1=L(M+1)$ for some positive integer $M+1$ and we can construct any such ruler from $L$ base rulers of order $M+1$ given by
$d^{(\ell)}_{0}<d^{(\ell)}_{1}<\dots<d^{(\ell)}_{M}$
for each $\ell\in[L]$ as
$\{d_{k}\mid k\in[M^{\prime}+1]\}$ =$=\{Ld^{(\ell)}_{k}+\ell\mid k\in[M+1],\ell\in[L]\}$
$=\bigcup_{\ell=0}^{L-1}\{Ld^{(\ell)}_{k}+\ell\mid k\in[M+1]\}$ (16)
$=\bigcup_{k=0}^{M}\{Ld^{(\ell)}_{k}+\ell\mid\ell\in[L]\}\text{.}$ (17)
Observe that (16) corresponds to a partition of the ruler into parts which are congruent modulo $L$ while (17) corresponds to a partition into parts which are not congruent modulo $L$. From (15) and (16), we see that an $L$-uniform ruler of order $L(M+1)$ leads to
@@
for all $(n,i,j)\in\mathcal{P}$ as needed.
@@ IV-B Scattering Conditions
Next, we observe that the inter-block scattering condition on (9) becomes that for fixed $n^{\prime}_{1},n^{\prime}_{2}\in L\mathbb{Z}$ with $n^{\prime}_{2}>n^{\prime}_{1}$,
@@
should have a unique solution $k^{\prime}_{1},k^{\prime}_{2}\in[M^{\prime}+1]$ if any which necessarily satisfies $k_{2}^{\prime}>k^{\prime}_{1}$ and $d_{k^{\prime}_{1}}\equiv d_{k^{\prime}_{2}}\pmod{*}{L}$ since $n^{\prime}_{2}-n^{\prime}_{1}\equiv 0\pmod{*}{L}$. Using an $L$-uniform ruler as in Definition 7, we see that we must have for some $\ell\in[L]$ that $d_{k^{\prime}_{1}},d_{k^{\prime}_{2}}\in\{Ld^{(\ell)}_{k}+\ell\mid k\in[M+1]\}$ thus
@@ IV-C Encoding
Next, observe that all base rulers can be translated so that $d_{0}^{(\ell)}=0$ for all $\ell\in[L]$. This results in
$(d_{0},d_{1},\dots,d_{L-1})=(0,1,\dots,L-1)$
for our $L$-uniform ruler $d_{0}<d_{1}<\dots<d_{M^{\prime}}$. This also leads to
@@ IV-D Higher-Order Staircase Codes
It remains to define difference triangle sets. In particular, an inter-block scattering $L$-uniform ruler of order $L(M+1)$ arises from an $(L,M)$-DTS defined as follows:
@@
An $(L,M)$-DTS is a set of $L$ Golomb rulers of order $M+1$ whose respective sets of distances are disjoint. Equivalently, it is a set of $L$ rulers of order $M+1$ given by
$d^{(\ell)}_{0}<d^{(\ell)}_{1}<\dots<d^{(\ell)}_{M}$
for $\ell\in[L]$ such that all positive differences
$d^{(\ell)}_{k_{2}}-d^{(\ell)}_{k_{1}}$
for $k_{1},k_{2}\in[M+1]$, $k_{2}>k_{1}$, and $\ell\in[L]$ are distinct.
Evidently, an $L$-uniform ruler of order $M^{\prime}+1=L(M+1)$ whose base rulers are an $(L,M)$-DTS leads to unique solutions to (18) thus inter-block scattering. The construction is summarized by the following definition, which we regard as the central definition, and main contribution, of this paper.
@@
$\pi^{\prime}_{2}$ $=\pi^{\prime}_{4}=\pi_{1}$
$\pi^{\prime}_{3}$ $=\pi^{\prime}_{5}=\pi_{2}\text{.}$
@@
Figure 2: Visualization of the code of Example 4; rows belong to $\mathcal{C}$.
A visualization of the code of Example 4 is provided in Fig. 2 in the style of zipper codes [3] which better illustrates the encoder structure.
Finally, we note that for higher-order staircase codes as just defined, the code rate, $4$-cycle-freeness, component code, and variable and constraint degrees are all independent of $L$. We therefore expect the threshold and error floor performance to be roughly independent of $L$. For the remainder of this paper, we will be interested in the effect of increasing $L$ while keeping all other parameters fixed.
IV-E Memory
@@
At first glance, encoding requires that memory of the last $d_{M^{\prime}}$ transmitted blocks be maintained. A naive choice would be to maintain the set of $L\cdot\max_{\ell\in[L]}d_{M}^{(\ell)}$ blocks $\mathsf{BUFF}_{\frac{n}{L}}$ defined by
@@
for encoding the $(n/L)$th information block. However, by looking at the ruler decomposition (16) we see that we can alternatively maintain $L$ independent buffers $\mathsf{BUFF}^{(\ell)}_{\frac{n}{L}}$ of sizes $d_{M}^{(\ell)}$ for $\ell\in[L]$ given by
@@
which leads to a potentially reduced memory requirement since
$\sum_{\ell=0}^{L-1}d_{M}^{(\ell)}\leq L\cdot\max_{\ell\in[L]}d_{M}^{(\ell)}\text{.}$ (21)
@@
bits (or nonbinary symbols).
@@
as a measure of the decoding memory.
We remark that while (23) is expected to be a reliable proxy for decoding memory assuming that the same trends illustrated in [3] hold for our more general code family, it is not true that (22) is necessarily a reliable proxy for encoding memory. This is because the entirety of every $(S/L)\times(S/L)$ block need not be stored. Instead, a pipelined approach can be employed in which blocks are compressed down to their respective future additive contributions to parity. An example of this can be found in [19] which provides such a hardware implementation of staircase encoders corresponding to $(L=1,M=1)$. However, even with such approaches, there can still be significant hardware efficiency benefits to increasing $L$ as demonstrated in [20] which compares a hardware encoder for a variant of tiled diagonal zipper codes corresponding to $(L=32,M=1)$ to the hardware encoder of [19].
IV-F Notions of Optimality for $(L,M)$-DTS Constructions

### v5

@@ IV-A Code Structure
Consider starting with a weak generalized staircase code of sidelength $S^{\prime}$ and degree parameter $M^{\prime}$. We further have our permutations of $[S^{\prime}]\times[S^{\prime}]$ given by $\pi^{\prime}_{k}$ for $k\in[M^{\prime}+1]$ and the corresponding matrix-permuting functions given by $\Pi^{\prime}_{k}$ for $k\in[M^{\prime}+1]$. Finally, we have a ruler $(d_{0},d_{1},\dots,d_{M^{\prime}})$ and a component code $\mathcal{C}$ of length $(M^{\prime}+1)S^{\prime}$ and dimension $(M^{\prime}+1)S^{\prime}-r$. However, we have the following twist: We define our constraint that the rows of
@@
belong to $\mathcal{C}$ as holding for all integers $n\in L\mathbb{Z}$ where $L$ is a new fixed positive integer parameter. In other words, we restrict the constraint on the rows of (14) to integers $n$ that are multiples of $L$.
Revisiting the Tanner graph perspective, the constraint nodes become $\mathcal{L}=L\mathbb{Z}\times[S^{\prime}]$ and from (5) we see that a variable node $(n,i,j)\in\mathcal{P}$ has the constraint $(n^{\prime},i^{\prime})\in\mathcal{L}$ as a neighbour if and only if $n^{\prime}=n+d_{k}$ for some $k\in[M^{\prime}+1]$ where $n^{\prime}\in L\mathbb{Z}$. Therefore, we get a variable node degree of
@@
which depends on the block index $n$ and the ruler. We seek to recover our original regular variable degree of $M+1$ for all variable nodes independently of $n$. This requires an $L$-uniform ruler defined as follows:
@@
A ruler $(d_{0},d_{1},\dots,d_{M^{\prime}})$ of order $M^{\prime}+1$ is $L$-uniform if it has an equal number of marks belonging to each of the residue classes
$L\mathbb{Z},L\mathbb{Z}+1,\dots,L\mathbb{Z}+L-1\text{.}$
Necessarily, any such ruler can be constructed from a set of $L$ base rulers $\mathcal{X}=\{X_{0},X_{1},\dots,X_{L-1}\}$ each of order $M+1=(M^{\prime}+1)/L$ as the union of the entries of
$LX_{0},LX_{1}+1,\dots,LX_{L-1}+L-1$
together with some ordering.
Remark 4.
Normalizing an $L$-uniform ruler produces a ruler which is still $L$-uniform.
Remark 5.
Given $L$ order $M+1$ (base) rulers
$X_{\ell}=(d_{0}^{(\ell)},d_{1}^{(\ell)},\dots,d_{M}^{(\ell)})$
for $\ell\in[L]$, the elements of the corresponding $L$-uniform ruler of order $L(M+1)$ can be arranged into an $L\times(M+1)$ array
$\begin{array}[]{llll}Ld_{0}^{(0)}&Ld_{1}^{(0)}&\cdots&Ld_{M}^{(0)}\\ Ld_{0}^{(1)}+1&Ld_{1}^{(1)}+1&\cdots&Ld_{M}^{(1)}+1\\ \phantom{L}\vdots&\phantom{L}\vdots&\ddots&\phantom{L}\vdots\\ Ld_{0}^{(L-1)}\!+\!L\!-\!1&Ld_{1}^{(L-1)}\!+\!L\!-\!1&\cdots&Ld_{M}^{(L-1)}\!+\!L\!-\!1\\ \end{array}$
where the rows partition it into parts that are congruent modulo L and the columns partition it into parts that are not congruent modulo L.
From (15) and Definition 8, we see that an $L$-uniform ruler of order $L(M+1)$ leads to
@@
for all $(n,i,j)\in\mathcal{P}$ as needed. Therefore, we can henceforth assume an $L$-uniform ruler as in Remark 5. We assume without loss of generality that both the base rulers and the corresponding $L$-uniform ruler are normalized. Before describing the encoding process for the family of codes obtained so far, we derive the scattering conditions which further reveal a simplification: Only $M+1$ distinct permutations are needed, one for each of the columns of the array in Remark 5.
@@ IV-B Scattering Conditions
Observe that the inter-block scattering condition on (9) becomes that for fixed $n^{\prime}_{1},n^{\prime}_{2}\in L\mathbb{Z}$ with $n^{\prime}_{2}>n^{\prime}_{1}$,
@@
should have a unique solution $k^{\prime}_{1},k^{\prime}_{2}\in[M^{\prime}+1]$ if any which necessarily satisfies $k_{2}^{\prime}>k^{\prime}_{1}$ and $d_{k^{\prime}_{1}}\equiv d_{k^{\prime}_{2}}\pmod{*}{L}$ since $n^{\prime}_{2}-n^{\prime}_{1}\equiv 0\pmod{*}{L}$. We must then have for some $\ell\in[L]$ that $d_{k^{\prime}_{1}},d_{k^{\prime}_{2}}\in\{Ld^{(\ell)}_{k}+\ell\mid k\in[M+1]\}$ thus
@@ IV-C Encoding
Note that if the base rulers are normalized, we get that $d_{0}^{(\ell)}=0$ for each $\ell\in[L]$. Observe then from Remark 5 that choosing the natural ordering for the corresponding $L$-uniform ruler so that $d_{0}<d_{1}<\cdots<d_{L(M+1)-1}$ not only results in a normalized ruler, but also results in
$(d_{0},d_{1},\dots,d_{L-1})=(0,1,\dots,L-1)\text{.}$
This leads to
@@ IV-D Higher-Order Staircase Codes
It remains to define difference triangle sets (DTSs):
@@
An $(L,M)$-DTS is set of $L$ Golomb rulers $\mathcal{X}=\{X_{0},X_{1},\dots,X_{L-1}\}$ each of order $M+1$ whose respective sets of distances are disjoint. Equivalently, $L$ order $M+1$ rulers given by
$X_{\ell}=(d_{0}^{(\ell)},d_{1}^{(\ell)},\dots,d_{M}^{(\ell)})$
where $\ell\in[L]$ form an $(L,M)$-DTS if all differences $d^{(\ell)}_{k_{1}}-d^{(\ell)}_{k_{2}}$ for $k_{1},k_{2}\in[M+1]$ with $k_{1}\neq k_{2}$ and $\ell\in[L]$, are distinct.
It then follows that an $L$-uniform ruler of order $M^{\prime}+1=L(M+1)$ constructed by taking the base rulers to be an $(L,M)$-DTS leads to unique solutions to (16) thus inter-block scattering. The construction is summarized by the following definition:
@@
Importantly, Section IV-B implies the following:
Proposition 3.
Higher-order staircase codes are scattering.
@@
$\pi^{\prime}_{4}$ $=\pi^{\prime}_{2}=\pi_{1}$
$\pi^{\prime}_{5}$ $=\pi^{\prime}_{3}=\pi_{2}\text{.}$
@@
Figure 4: Visualization of a tiled diagonal zipper code corresponding to $L=2$, $M=1$, $(d_{1}^{(0)},d_{1}^{(1)})=(1,2)$, and $\Pi_{1}(B)=B^{\mathsf{T}}$; rows belong to $\mathcal{C}$.
Figure 5: Visualization of the code of Example 4 where $L=2$ and $M=2$; the respective sets of distances of the two Golomb rulers are disjoint and rows belong to $\mathcal{C}$.
A visualization of an $(L=2,M=1)$ code corresponding to an instance of tiled diagonal zipper codes [11] is provided in Fig. 4. A visualization of the $(L=2,M=2)$ code of Example 4 follows it in Fig. 5 illustrating an increase in variable degree while maintaining scattering. In both of these pictures, the $L=2$ rightmost columns represent the blocks that are actually transmitted while the columns to their left are delayed and permuted copies that are used to constrain the transmitted blocks.
Finally, we note that for higher-order staircase codes as just defined, the code rate, $4$-cycle-freeness, component code, and variable and constraint node degrees are all independent of $L$. It is therefore natural to study the effect of increasing $L$ while keeping all else fixed. We will find that while increasing $L$ reduces the block size to $1/L^{2}$ times its size, the number of blocks that must be kept in memory is $\Omega(L^{2})$ meaning that we can at most reduce the memory by a constant factor. However, this factor can still be meaningful in practice and we may nonetheless need to minimize memory for any given choice of $L$. Sections VI and VII will deal with exhibiting DTSs that lead to memory-optimal realizations of higher-order staircase codes.
Before proceeding, we comment briefly on the practical code performance implications of increasing $L$. Since the code rate, degree distribution, $4$-cycle-freeness, and component code are all unaffected by $L$, coding folklore suggests that the threshold and error floor performance should be roughly independent of $L$. For example, in the case of tiled diagonal zipper codes corresponding to $(L\geq 1,M=1)$, it is found in [3] that the gap to the hard-decision Shannon limit is increased by no more than $0.04$ dB as $L$ is increased from $L=1$ to $L=S=1000$ with the component code being a shortened triple-error-correcting BCH code of length $2000$ and dimension $1967$. A similar invariance can be observed for general higher-order staircase codes by experimenting with our simulators for these codes available available online [21, 25]. The reader is also referred to [26] for recent work on theoretical performance analysis under similar code families and decoding methods to what is considered in this paper. Moreover, the bound of Proposition 1 is independent of $L$ providing the same guarantee on the coarse asymptotic behaviour of the error floor.
IV-E Multiply-Chained and Other Extensions
We now adapt a technique of Zhao et al. [13] in which multiple tiled diagonal zipper codes [11] are chained in a circle to boost their performance. This is straightforwardly applicable to higher-order staircase codes. We introduce a new positive integer parameter $C>1$ and consider $C$ higher-order staircase code encoders indexed by $c\in[C]$ and acting on $C$ independent input streams with the following twist: The $c$th encoder uses the memory of the $(c-1\pmod{*}{C})$th encoder. This leads to multiply-chained higher-order staircase codes formally defined as follows:
Definition 11 (Multiply-chained higher-order staircase code).
Given all four of the following:
  • an integer $C>1$;
  • an $(L,M)$-DTS given by
    $0=d^{(\ell)}_{0}<d^{(\ell)}_{1}<\dots<d^{(\ell)}_{M}$
    for $\ell\in[L]$ with corresponding $L$-uniform ruler of order $L(M+1)$
    $d_{0}<d_{1}<\cdots<d_{L(M+1)-1}$
    given accordingly as
  • a component code $\mathcal{C}$ of length $(M+1)S$ and dimension $(M+1)S-r$,
a multiply-chained higher-order staircase code of rate $1-r/S$ is defined by the constraint on the $C$ bi-infinite sequences of $(S/L)\times(S/L)$ matrices $\dots,B_{-2}^{(c)},B_{-1}^{(c)},B_{0}^{(c)},B_{1}^{(c)},B_{2}^{(c)},\dots$ indexed by $c\in[C]$ that the rows of
$\bigg{(}\Pi^{\prime}_{L(M+1)-1}(B_{n-d_{L(M+1)-1}}^{(c-1\pmod{*}{C})})\bigg{|}\\ \cdots\bigg{|}\Pi^{\prime}_{L}(B_{n-d_{L}}^{(c-1\pmod{*}{C})})\bigg{|}\Pi^{\prime}_{L-1}(B_{n-d_{L-1}}^{(c)})\bigg{|}\cdots\bigg{|}\Pi^{\prime}_{0}(B_{n-d_{0}}^{(c)})\bigg{)}$
for each $c\in[C]$ belong to $\mathcal{C}$ for all $n\in L\mathbb{Z}$.
We define a multiply-chained higher-order staircase code with $C=1$ to simply be a higher-order staircase code as in Definition 10 so that we have a unified family defined for all $C\geq 1$. Lastly, we include one further generalization of this family which is as follows: Rather than considering one component code $\mathcal{C}$, we can consider a “time-varying” family of component codes such as $\{\mathcal{C}_{i}^{(c)}\mid c\in[C],i\in[S/L]\}$ where the $i$th row in the constraint span in Definition 11 belongs to $\mathcal{C}_{i}^{(c)}$. For example, this family can simply consist of $S/L$ permuted copies of a fixed code $\mathcal{C}$. With this extension, the family of Definition 11 includes the OFEC code family of [12] as the special case of $C=2$, $M=1$, and a special choice of time-varying component code permutation defined therein.
IV-F Memory
@@
We now consider the encoding memory requirements for higher-order staircase codes. For the multiply-chained variation, all requirements are simply multiplied by $C$ so we restrict our attention to the case of $C=1$. At first glance, encoding requires that memory of the last $d_{M^{\prime}}$ transmitted blocks be maintained. A naive choice would be to maintain the set of $L\cdot\max_{\ell\in[L]}d_{M}^{(\ell)}$ previous blocks $\mathsf{BUFF}_{\frac{n}{L}}$ defined by
@@
for encoding the $(n/L)$th information block. However, we can observe in the case of Example 4 illustrated in Fig. 5 that blocks with even indices must be recalled for longer than blocks with odd indices. In general, by examining the ruler decomposition given in Remark 5, we see that we can alternatively maintain $L$ independent buffers $\mathsf{BUFF}^{(\ell)}_{\frac{n}{L}}$ of sizes $d_{M}^{(\ell)}$ for $\ell\in[L]$ given by
@@
so that the total number of blocks stored is
$\sum_{\ell=0}^{L-1}d_{M}^{(\ell)}\leq L\cdot\max_{\ell\in[L]}d_{M}^{(\ell)}\text{.}$
@@
bits (or nonbinary symbols). This is analogous to the concept of total memory or overall constraint length in the theory of convolutional codes [27, Ch. 2] with the key differences being that bits are replaced with $(S/L)\times(S/L)$ blocks which admit permutation operations and that one-bit memory elements are replaced accordingly with stores of $(S/L)^{2}$ bits. The encoder style assumed here is analogous to Forney’s obvious realization [28].
@@
as a measure of the decoding memory. This is analogous to the constraint length in the theory of convolutional codes.
We emphasize that (19) and (20) are merely models for the encoding and decoding memory under a certain encoder and decoder structure considered in [11] which can be subverted in practice. There are two main ways in which this is done:
  • Each $(S/L)\times(S/L)$ block is compressed down to a $(S/L)\times r$ future additive contribution to parity and thus is not directly stored.
  • An encoder style analogous to the observer canonical form as opposed to the controller canonical form [27, Ch. 2] or obvious realization [28] is used in which the encoding memory is proportional to $\max_{\ell\in[L]}d_{M}^{(\ell)}$ rather than $\sum_{\ell=0}^{L-1}d_{M}^{(\ell)}$. This encoder style was first recognized by Massey [29].
Both of these are done in the hardware implementations of a staircase code encoder corresponding to $(L=1,M=1)$ and a zipper code encoder corresponding to $(L=32,M=1)$ found in [30] and [31] respectively. It is found in [31] that there are significant hardware efficiency benefits to $L=32$ relative to $L=1$. Nonetheless, one or both of $\sum_{\ell=0}^{L-1}d_{M}^{(\ell)}$ and $\max_{\ell\in[L]}d_{M}^{(\ell)}$ are measures of hardware complexity. Thus, we seek, for a given $L$ and $M$, an $(L,M)$-DTS for which $\sum_{\ell=0}^{L-1}d_{M}^{(\ell)}$ and/or $\max_{\ell\in[L]}d_{M}^{(\ell)}$ are minimized. We defer the study of this problem to Section VI.
V Extended-Hamming-Based Higher-Order Staircase Codes
We consider now the use of memory-optimal higher-order staircase codes and their multiply-chained variation with extended Hamming component codes. As discussed in the previous section, the construction of such codes requires the construction of DTSs with minimum scope and sum-of-lengths. Such DTSs are provided in the Appendix with the subject of how they are obtained deferred to Sections VI and VII. We use the optimal DTSs from the Appendix together with the extended Hamming codes defined by Table V to construct our codes.
The encoder produces a $C(S/L)\times L(S/L)$ rectangle of coded bits since each of the $C$ higher-order staircase code encoders produces $L$ blocks at a time each of dimensions $(S/L)\times(S/L)$. The decoding window size $W$ counts the number of such rectangles meaning that the number of bits in the decoding window is $WC(S/L)^{2}L$. Similarly to before, we assume transmission across a BSC and perform sliding window decoding of $W$ consecutive rectangles at a time where one iteration comprises decoding all constraints in the window consecutively and $I$ iterations are performed before advancing the window by one rectangle. We again use a termination-like procedure as in Section III in which a frame is formed from $F$ consecutive rectangles. We take $F=10^{5}+W$ for all simulations in this section so that the entire coding scheme is determined by the six-tuple of parameters $(L,M,S/L,C,W,I)$. The simulation curves in Figures 6 and 7 are labeled with this six-tuple appended with the resulting unterminated code rate $R_{\mathsf{unterminated}}$ and the decoding window size in bits for the reader’s convenience.
Figure 6: Simulation results for $R_{\mathsf{unterminated}}\approx 0.937$ examples with parameters $(L,M,S/L,C,W,I,R_{\mathsf{unterminated}},WC(S/L)^{2}L)$
Figure 7: Simulation results for $R_{\mathsf{unterminated}}\approx 0.87$ examples with parameters $(L,M,S/L,C,W,I,R_{\mathsf{unterminated}},WC(S/L)^{2}L)$
One of our goals in this section is to demonstrate the ability to construct a variety of codes with similar properties by varying $L$ at a roughly constant code rate. We accordingly organize our simulations slightly differently with results for two different approximate code rates given in Figures 6 and 7 respectively. The dashed lines indicate points at which zero bit errors were measured after transmitting a number of bits equal to ten times the reciprocal of the corresponding output BER.
Observe firstly that Fig. 7 illustrates control over the error floor and threshold via $M$ and $C$ respectively as predicted. Moreover, the examples in these figures perform comparably to the codes compared in [2, 13] while using only Hamming components as opposed to BCH components and also improve over the results of Section III with generalized staircase codes $(L=1)$ in terms of memory and decoding iterations.
We provide heuristic estimates of the performance, latency, and complexity for some examples as before. In Table VII, we compare the classical staircase code of [2] to the extended-Hamming-based higher-order staircase code with parameters $(L,M,S/L,C,W,I)=(7,4,25,1,162,1)$. As before, we assume that for the relatively small values of $M$ considered, the cost of each component code decoding is dominated by the cost of invoking the component code decoder $t^{2}$ rather than the cost of updating the syndromes of codewords involved in each corrected bit. In Table VIII, we compare a multiply-chained tiled diagonal zipper code from [13] with parameters $(L,M,S/L,C,W,I)=(2,1,125,2,8,4)$ and BCH component codes with $t=3$ to the extended-Hamming-based higher-order staircase code with parameters $(L,M,S/L,C,W,I)=(4,4,19,1,96,1)$. In both situations, we see large reductions in latency and complexity at a performance cost of roughly $0.3$ to $0.5$ dB.
TABLE VII: Heuristic estimates of performance, latency, and complexity for higher-order staircase codes with $R_{\mathsf{unterminated}}\approx 0.937$
staircase [2, 22] this work
gap to Shannon at 1e-15 0.56 dB 0.89 dB
latency: $WC(S/L)^{2}L$ 1.8e6 7.1e5
iterations: $I$ 3 1
decodings per iteration: $WC(S/L)$ 3.6e3 4.1e3
cost per decoding: $t^{2}$ 9 1
complexity score: $IWC(S/L)t^{2}$ 9.7e4 4.1e3
TABLE VIII: Heuristic estimates of performance, latency, and complexity for higher-order staircase codes with $R_{\mathsf{unterminated}}\approx 0.87$
Fig. 13 in [13] this work
gap to Shannon at 1e-8 0.75 dB 1.25 dB
latency: $WC(S/L)^{2}L$ 5e5 1.4e5
iterations: $I$ 4 1
decodings per iteration: $WC(S/L)$ 2e3 1.8e3
cost per decoding: $t^{2}$ 9 1
complexity score: $IWC(S/L)t^{2}$ 7.2e4 1.8e3
We anticipate improved performance–complexity–latency tradeoffs in a wider variety of contexts in light of recent works [32, 33] which accomplish such goals via new product-like codes with bit degree increased from two to three. The simulations here are only meant to provide a few motivating examples. These results can be reproduced and extended using our simulation tools available online [21, 25].
VI Difference Triangle Sets with Minimum Scope and Sum-of-Lengths
Consider an $(L,M)$-DTS given by
$0=d^{(\ell)}_{0}<d^{(\ell)}_{1}<\dots<d^{(\ell)}_{M}$